-- PaulSmith - 22 Feb 2007

SDT Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith

Inst. Support: Olga Kuhn TO: Ed Lundin

1835: Flats UBVgr (0 deg PA); telescope not tracking very well. Bright stars seen in flats are short streaks. Only 3 exposures done with r filter.

1908: Move to pointing stars to see if telescope is pointing where it thinks it is pointing to try to resolve tracking issue during flat fields. Pointing is way off. Work had been done on the encoders during the day.

2015: Try pointing star to see if the telescope has found itself again. It has. Go off to standard star field for a first focus. OB=SA98_BV; seeing.

2025: Off to KBO1 and do WF. Then start first sequence of OBs at 2032. seeing 0.8-1.1". Last exposure in sequence started 2110.

2115: Move to M67, do WF, then OB=WDs_M67_SCI_V3; lost focus, during OB. Abort OB then WF then redo OB. Same image quality problems, so we decide to abort and go back to KBO1.

2145: Back to KBO1 and do WF. 2150 do first exposure of second KBO1 pass, but can't get an in-focus image. Reset primary and do more WFs. Trouble getting coma to settle down. Consult John Hill on optical issues.

[JMH: After studying the images more carefully during the day, I think the main problem is still field dependent coma from a cause still unknown to us (see Issue #506). Tonight a bit of astigmatism and seeing mixed with the coma makes it less obvious than it was in the superb conditions of last night.]

2212: Start second pass of KBO1. Image quality not so great, but need to press forward. Now we have scattered light problems. More delays as we sort through various issues. Take astigmatism corrections out. Images better so we continue with KBO1; seeing better than 1.0". Scattered light problem seems to be a one-time thing (not moon). Maybe we got hit with a headlight. Will redo that field. Seeing better than 1" in most frames, however there is still some significant variations in PSF across the array, though this is not nearly as bad as when the problem cropped up after going to M67 and the science should not be compromised. KBO1 finally completed at 2335. Whew!!

[JMH: The "scattered light" image 052904 is quite interesting. I believe that an airplane flew across your field (note diagonal stripe) and flashed one of his wing strobes while in the middle of chip4 (note 1000 pixel diameter pupil). Based on the diameter of the pupil, the plane was at an altitude of 7 km above the telescope.]

2340 go to COSMOS field, do WF, then OB=cosmos_rSLOAN_1; seeing around 1". Moon has set.

0020: WF and then OB=cosmos_rSLOAN_2; seeing 1.0-1.4".

0100: Move to Lockman hole and do WF. OB=Egami_Lockman_VLA_B_2; seeing 1.3-1.4", not quite what program needs. Finish OB and then head for a target more tolerant of worsening seeing. Still clear, but California front is approaching.

0145: Try to move but the telescope doesn't want to move. Possible trouble with frozen LBC window. Reboot LBC computer to try to get things started again.

0218: Finally get LBC control up and running again. No known reason for the freeze up and no known reason why is decided to come back (loss of nearly an hour). Do WF, then OB=fuobj1_1_g; seeing 1.1-1.3". Wind picking up as front nears.

0300: Go to first Quasar field, WF, then OB=sdt_fan1_1; seeing (at U) 1.7"; abort OB since we didn't want to do U imaging in these conditions. Will try different filter. Smith highly recommends that PIs place some indication of the filter to be used in the OB title, not buried in some README file. Having to carefuly read filter coding after midnight on a night with lots of technical troubles is asking too much from simple queue observers. Start OB=sdt_fan_1_1 (SLOAN r and g) at 0315; seeing 1.2-1.6". Do WF but pupils are so bad we can't get a sensible solution.

0353: Go to standard star field to get more stars for the WF. Dead in the water. More issues arise concerning the operation of the LBC gui and its communication with TCS and primary mirror comtrol. Olga trying hard to sort this out, but the problem(s) seems to be quite complex and effectively the night is over.
Topic revision: r5 - 25 Feb 2007, PaulSmith
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback